OpenStudy (anonymous):

How would you Simplify 4 x all over 9 plus quantity x minus 10 over 9?????

OpenStudy (anonymous):

\[\frac{4x}{9}-\frac{x-10}{9}\] \[=\frac{4x-(x-10)}{9}\] is a start

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

:/ I wish some one helped me.